# Homogeneous Tokenizer Matters: Homogeneous Visual Tokenizer for Remote Sensing Image Understanding

### v1

@@ Abstract
The paradigm shift introduced by multimodal large language models, which is based on the transformer architecture and the pretext task of ”next-token prediction,” has revolutionized the field of remote sensing image understanding. However, the tokenizer, as one of the fundamental components of large models, has long been overlooked or even misunderstood in visual tasks. One key factor of the great comprehension power of the large language model is that natural language tokenizers utilize meaningful words or subwords as the basic elements of language. In contrast, mainstream visual tokenizers, represented by patch-based methods such as Patch Embed, rely on meaningless rectangular patches as basic elements of vision, which cannot serve as effectively as words or subwords in language. Starting from the essence of the tokenizer, we defined semantically independent regions (SIRs) for vision. Based on this definition, we propose two properties that an ideal visual tokenizer should possess: (1) homogeneity, where SIRs serve as the basic elements of vision, and (2) adaptivity, which allows for a flexible number of tokens to accommodate images of any size and tasks of any granularity. Based on this, we designed a simple HOmogeneous visual tOKenizer: HOOK. HOOK mainly consists of two modules: the Object Perception Module (OPM) and the Object Vectorization Module (OVM). To achieve homogeneity, the OPM splits the image into 4×4 pixel seeds and then utilizes the attention mechanism to perceive SIRs. The OVM employs cross-attention to merge seeds within the same SIR. To achieve adaptability, the OVM defines a variable number of learnable vectors as cross-attention queries, allowing for the adjustment of token quantity. We conducted experiments on the NWPU-RESISC45, WHU-RS19 classification dataset, and GID5 segmentation dataset for sparse and dense tasks. The results demonstrate that the visual tokens obtained by HOOK correspond to individual objects, which demonstrates homogeneity. HOOK outperformed Patch Embed by 6% and 10% in the two tasks and achieved state-of-the-art performance compared to the baselines used for comparison. Compared to Patch Embed, which requires more than one hundred tokens for one image, HOOK requires only 6 and 8 tokens for sparse and dense tasks, respectively, resulting in efficiency improvements of 1.5 to 2.8 times. The homogeneity and adaptability we proposed provide new perspectives on the study of visual tokenizers, and guided by these properties, the HOOK we designed shows potential to replace Patch Embed. The code is available at https://github.com/GeoX-Lab/Hook.
@@
Remote Sensing Image Understanding , Visual Tokenizer , Homogeneous , Semantic Independent Region , Visual Transformer Model
@@
Figure 1: (a) Natural language tokenizer uses words or subwords as the basic elements of language. Similarly, for visual tokenizers, this work attempts to answer a fundamental question: What are the basic elements of vision? (b) The current mainstream visual tokenizers are patch-based methods, represented by Patch Embed, that can be categorized into three types based on hierarchy: patch level, subpatch level, and superpatch level. Their commonality lies in the patches as the basic elements of vision. (c) There exists a confusion matrix for tokens and objects. “Same Object Multiple Tokens” leads to incomplete object features, while “Same Token Multiple Objects” leads to unclear relationships between objects. “Multiple Tokens Multiple Objects” inherits the drawbacks of the above two cases. Patch-based methods inherently struggle to achieve the ideal scenario: ”Same Token Same Object”. (d) We define the concept of the Semantic Independent Region (SIR) and propose two properties of an ideal visual tokenizer: homogeneity and adaptability. We design a simple HOmogeneous visual tOKenizer, HOOK, where SIRs as the basic elements of vision.
@@ 1 Introduction
Large language models, such as the GPT series[1, 2, 3, 4], Llama[5, 6] and PaLM[7] which are built on transformer architecture and trained with the ”next-token prediction” objective, have demonstrated powerful natural language understanding, reasoning, and generalization capabilities. Subsequently, the development of multimodal large language models[8, 9, 10, 11, 12] has thoroughly transformed the research landscape in various fields, including remote sensing image understanding, leading to a paradigmatic revolution.
@@
The tokenizer is a foundational and essential component of multimodal large language models that is aimed at identifying the basic elements of the data and using them to tokenize the data into a sequence of tokens. The natural language tokenizer (NL tokenizer) is fundamental for machine understanding of language and has been widely investigated and applied in the field of natural language processing[13, 14, 15, 16]. As shown in Figure 1-a, one key factor contributing to the great comprehension power of the large language model is that natural language tokenizers utilize meaningful words or subwords as the basic elements of language. However, visual tokenizers, which play a similar role in visual foundation models and multimodal large models, have long been overlooked, underestimated, and even misunderstood.
@@
The visual tokenizer is aimed at identifying the basic elements of an image and tokenizing the image into a sequence of tokens based on the basic elements. What are the basic elements of vision? As shown in Figure 1-b, patch-based methods, represented by Patch Embed[17], are the most popular implementations of visual tokenizers, where the fundamental characteristic of these methods is the use of rectangular patches as the basic elements of vision. The advantage of patch-based methods lies in their simplicity and efficiency, but they have two prominent issues for remote sensing images:
@@
(1) Rectangular patches fail to match irregular and complex objects, which leads to a lack of semantic homogeneity within tokens. As shown in Figure 1-c, there exists a confusion matrix for tokens and objects: if multiple objects are aggregated within a single token, the model may struggle to learn the relationships between the objects; if a single object is dispersed across multiple tokens, the model may struggle to learn the complete features of the individual object. We summarize these two scenarios as follows: “Same Token Multiple Objects” and “Same Object Multiple Tokens”. Additionally, if multiple objects correspond to multiple tokens, i.e., “Multiple Tokens Multiple Objects”, it inherits the drawbacks of the previous scenarios. Ideally, the relationship between tokens and objects should be “Same Object Same Token”, but due to the limitations of rectangular patches, patch-based methods fundamentally cannot achieve this ideal scenario.
@@
(2) The efficiency issues arising from the fixed and redundant number of tokens limit the application of transformer-based models in remote sensing images. For instance, when the patch size is 16 $\times$ 16, a 224 $\times$ 224 remote sensing image corresponds to 196 tokens, and the number of tokens grows quadratically with increasing image size. Furthermore, in transformer models, the computational complexity of the attention mechanism is $O(N^{2})$, which indicates that the computational cost of the attention mechanism also scales quadratically with the number of tokens. This feature makes it challenging for transformer-based remote sensing models to handle high-resolution remote sensing images.
@@
Therefore, meaningless rectangular patches are not suitable as the basic elements of images. Analogous to meaningful words or subwords that serve as the basic elements of natural language, an intuitive idea is to identify certain semantically meaningful regions in the image as the basic elements of the vision. As shown in Figure 1-d, we expand this idea and define two fundamental properties that an ideal visual tokenizer should possess:
(1) Homogeneity: Inspired by the NL tokenizer, which uses meaningful words or subwords as the basic elements of natural language, we define semantically independent regions (SIRs) for images. SIRs refer to regions in the image that are semantically independent of other regions outside themselves. Once SIRs are defined, they are internally homogeneous in semantics compared to inter-regions; thus, we refer to the visual tokenizer with SIRs as the basic elements of vision as homogeneous.
(2) Adaptivity: In the definition of SIRs, ”independence” is a dynamic concept. For example, the entire aircraft is an independent region compared to the ground, the wing is an independent region compared to the aircraft, and the flap is an independent region compared to the wing. Therefore, when addressing a specific remote sensing image, the image size and the granularity of the task determine the ideal granularity of independence, while the number of tokens determines the actual granularity of independence. Additionally, a visual tokenizer should be image-agnostic and task-agnostic. Therefore, the number of tokens should be adjusted to accommodate images of any size and tasks of any granularity. We refer to this property as the adaptivity of the visual tokenizer.
@@
To develop an ideal visual tokenizer that meets the above properties, we define the confusion matrix for tokens and objects under strict definitions and constraints. We found that the transformations among “Same Object Same Token”, “Same Object Multiple Tokens”, “Same Token Multiple Objects”, and “Multiple Tokens Multiple Objects” actually involve only two meta-operations: ”Split” and ”Merge”. The process of obtaining homogeneous visual tokens can be abstracted as a routing selection problem, where two distinct general routings are ”Splitting and Merging” and ”Merging and Splitting”.
@@
Based on the ”Splitting and Merging” routing, we designed a simple HOmogeneous visual tOKenizer: HOOK. HOOK mainly consists of two modules: the object perception module (OPM) and the object vectorization module (OVM).
@@
To achieve homogeneity, HOOK adopts a strategy of first splitting the image into the finest granularity possible and then gradually merging to obtain SIRs. First, the OPM first leverages convolutional blocks to obtain 4 $\times$ 4 pixel seeds. Second, self-attention layers are used to associate the seeds belonging to the same SIR. Last, the OVM employs cross-attention to merge the associated seeds to obtain homogeneous visual tokens.
@@
To achieve adaptivity, the OVM defines $N$ learnable vectors as queries for the cross-attention module, with the seeds outputted by the OPM serving as the key and value. Here, $N$ is treated as a variable hyperparameter to allow for the arbitrary adjustment of token quantity.
@@
We tested the performance of HOOK on the sparse task using the NWPU-RESISC45[18] and WHU-RS19[19, 20] classification datasets and on the dense task using the GID5[21] semantic segmentation dataset. The experimental results show that the visual tokens obtained by HOOK correspond to individual objects, which demonstrates homogeneity. Compared to Patch Embed, HOOK achieved an improvement greater than 6% and 10% on the sparse task and dense task, respectively. Additionally, HOOK outperformed the baseline methods that we used for comparison, reaching a state-of-the-art performance. In terms of efficiency, HOOK requires only 6 tokens for the sparse task and 8 tokens for the dense task, whereas Patch Embed requires hundreds of tokens for a single image. This overall efficiency improvement entails an increase of 1.5 to 2.8 times.
@@
(1) Our analyses and experiments show that the importance of visual tokenisation is far underestimated. Starting from the essence of tokenizers, we propose two fundamental properties that an ideal visual tokenizer should possess: homogeneity and adaptivity.
(2) To develop a visual tokenizer that satisfies homogeneity and adaptivity, we rigorously define the confusion matrix for tokens and objects under strict definitions and constraints. We argue that the construction of homogeneous tokens can be abstracted as a routing selection problem, leading to two general routings: ”Splitting and Merging” and ”Merging and Splitting”. This analysis provides new perspectives and ideas for research on visual tokenizers.
(3) Based on the ”Splitting and Merging” routing, we designed a simple homogeneous visual tokenizer, HOOK, which offers an object-oriented, plug-and-play implementation for visual tokenizers.
(4) The experimental results demonstrate that HOOK satisfies homogeneity and adaptivity. HOOK outperforms baselines, including Patch Embed, in terms of both performance and efficiency. This finding shows the potential for HOOK to replace patch-based methods and become a new foundation visual tokenizer.
We organize this paper as follows: In Section 2, we review the advancements in visual tokenizers, which can be classified into two types—patch-based and object-oriented on the different basic elements of vision. In Section 3, we conduct a theoretical analysis of visual tokens. We start by emphasizing the importance and necessity of homogeneity from the essence of visual tokens. We then rigorously define and analyze the confusion matrix for tokens and objects to identify two general routings for constructing homogeneous visual tokens. Sections 4 and 5 introduce HOOK, a homogeneous visual tokenizer designed based on the ”Splitting and Merging” routing, and discuss its performance via experiments. The results indicate that HOOK satisfies both homogeneity and adaptivity. Section 6 primarily delves into the significance of visual tokenizers, rethinks our HOOK, and outlines future research directions.
@@ 2 Related Work
2.1 Patch-based Visual Tokenizer
Patch-based methods use rectangular patches as the basic elements of images. Patch Embed, which is widely applied in patch-based methods, splits an image into nonoverlapping patches of a fixed size, typically 16×16 pixels. The advantage of Patch Embed lies in its simplicity and efficiency. However, the redundant token quantity and fixed patch size have been identified in various works[22, 23, 24, 25] as factors of the instability of training in ViT models.
As shown in Figure 1-b, we categorize the improvements in patch-based methods into three types based on hierarchy: patch level, subpatch level, and superpatch level.
@@ A. Patch-level Methods
Patch-level methods argue that the visual tokens obtained through Patch Embed are highly redundant, and in reality, the model needs to select only a small portion of these tokens to recognize the image.
H. Yin et al. proposed A-ViT[26], where they defined the first dimension of tokens as the probability of discarding the token itself, which enables the model to autonomously select tokens that aid in image recognition. Y. Tang et al. calculated the impact of each token at each layer on the final model output, identifying and removing redundant tokens[27]. Y. Rao et al. introduced DynamicViT[28], and B. Pan et al. proposed IA-RED2[29], which inserted trainable prediction modules in the middle layers of the model for dynamic token selection. D. Marin et al. utilized pooling strategies to discard redundant tokens[30]. Y. Liu et al. introduced PatchDropout[31], a simple method that randomly drops a certain percentage of patches during model training, leading to efficiency improvements without compromising test accuracy. Y. Liang et al. proposed a token reorganization module to identify and merge tokens with minimal relevance to the CLS token, reducing the number of tokens[32]. Similarly, T. Wang et al. proposed PnP-DETR[33], which initially extracts features from the original image using ResNet as tokens and then identifies and merges background tokens, effectively enhancing model efficiency and object-level perception capabilities.

### v2

@@ Abstract
On the basis of the transformer architecture and the pretext task of ”next-token prediction”, multimodal large language models (MLLMs) are revolutionizing the paradigm in the field of remote sensing image understanding. However, the tokenizer, as one of the fundamental components of MLLMs, has long been overlooked or even misunderstood in visual tasks. A key factor contributing to the great comprehension power of large language models is that natural language tokenizers utilize meaningful words or subwords as the basic elements of language. In contrast, mainstream visual tokenizers, represented by patch-based methods such as Patch Embed, rely on meaningless rectangular patches as basic elements of vision. Analogous to words or subwords in language, we define semantically independent regions (SIRs) for vision and then propose two properties that an ideal visual tokenizer should possess: (1) homogeneity, where SIRs serve as the basic elements of vision, and (2) adaptivity, which allows for a flexible number of tokens to accommodate images of any size and tasks of any granularity. On this basis, we design a simple HOmogeneous visual tOKenizer: HOOK. HOOK consists of two modules: an object perception module (OPM) and an object vectorization module (OVM). To achieve homogeneity, the OPM splits the image into 4$\times$4 pixel seeds and then uses a self-attention mechanism to identify SIRs. The OVM employs cross-attention to merge seeds within the same SIR. To achieve adaptability, the OVM predefines a variable number of learnable vectors as cross-attention queries, allowing for the adjustment of the token quantity. We conducted experiments on the NWPU-RESISC45, WHU-RS19, and NaSC-TG2 classification datasets for sparse tasks and the GID5 and DGLCC segmentation datasets for dense tasks. The results show that the visual tokens obtained by HOOK correspond to individual objects, thereby verifying their homogeneity. Compared with randomly initialized or pretrained Patch Embed, which required more than one hundred tokens per image, HOOK required only 6 and 8 tokens for sparse and dense tasks, respectively, resulting in performance improvements of 2% to 10% and efficiency improvements of 1.5 to 2.8 times. The homogeneity and adaptability of the proposed approach provide new perspectives for the study of visual tokenizers. Guided by these principles, the developed HOOK has the potential to replace traditional Patch Embed. The code is available at https://github.com/GeoX-Lab/Hook.
@@
Remote Sensing Image Understanding , Visual Tokenizer , Homogeneous , Semantically Independent Region , Visual Transformer Model Citation: Run Shao, Zhaoyang Zhang, Chao Tao, Yunsheng Zhang, Chengli Peng, Haifeng Li. Homogeneous Tokenizer Matters: Homogeneous Visual Tokenizer for Remote Sensing Image Understanding. ISPRS Journal of Photogrammetry and Remote Sensing. 2024, 18: 294-310. DOI: 10.1016/j.isprsjprs.2024.09.009.
@@
Figure 1: (a) A natural language tokenizer uses words or subwords as the basic elements of language. Similarly, for visual tokenizers, this work attempts to answer a fundamental question: What are the basic elements of vision? (b) The current mainstream visual tokenizers are patch-based methods, represented by Patch Embed, which can be categorized into three types on the basis of hierarchy: patch level, subpatch level, and superpatch level. Their commonality lies in the use of patches as the basic elements of vision. (c) There exists a confusion matrix for tokens and objects. The ”same object multiple tokens” leads to incomplete object features, whereas the ”same token multiple objects” leads to unclear relationships between objects. The ”multiple tokens multiple objects” inherits the drawbacks of the above two approaches. Patch-based methods inherently struggle to achieve the ideal ”same token same object” scenario. (d) We define the concept of semantically independent region (SIRs) and propose two properties of an ideal visual tokenizer: homogeneity and adaptability. We design a simple HOmogeneous visual tOKenizer, HOOK, where SIRs as the basic elements of vision.
@@ 1 Introduction
Large language models, such as the GPT series[1, 2, 3, 4], Llama[5, 6], and PaLM[7], which are built on the transformer architecture and trained with the ”next-token prediction” objective, have demonstrated powerful natural language understanding, reasoning, and generalization capabilities. Subsequently, the development of multimodal large language models[8, 9, 10, 11, 12] has dramatically reshaped the research landscape in various fields, including remote sensing image understanding.
@@
Tokenizers are a foundational and essential component of multimodal large language models. They are designed to identify basic elements of input data and convert them into a sequence of tokens. Natural language tokenizers (NL tokenizers) are fundamental for machine understanding of language and have been extensively researched in the field of natural language processing[13, 14, 15, 16]. As shown in Figure 1-a, one key factor contributing to the great comprehension power of large language models is that natural language tokenizers utilize meaningful words or subwords as the basic elements of language. However, visual tokenizers, which play a similar role in visual foundation models and multimodal large models, have long been overlooked, underestimated, and even misunderstood.
@@
A visual tokenizer is aimed at identifying the basic elements of an image and tokenizing them into a sequence of tokens. What are the basic elements of vision? As shown in Figure 1-b, patch-based methods, represented by Patch Embed[17], are the most popular implementations of visual tokenizers. The fundamental characteristic of these methods is the use of rectangular patches as the basic elements of vision. The advantage of patch-based methods lies in their simplicity and efficiency, but they have two prominent issues for remote sensing images:
@@
(1) Rectangular patches fail to match irregular and complex objects. As shown in Figure 1-c, according to the confusion matrix for tokens and objects, if multiple objects are aggregated within a single token, the model may struggle to learn the relationships between the objects, whereas if a single object is dispersed across multiple tokens, the model may struggle to learn the complete features of the individual object. We refer to these two scenarios as the ”same token multiple objects” and ”same object multiple tokens” scenarios, respectively. Additionally, if multiple objects correspond to multiple tokens, i.e., ”multiple tokens multiple objects”, this scenario inherits the drawbacks of the previous scenarios. Ideally, the relationship between tokens and objects should be ”same object same token”, but owing to the limitations of rectangular patches, patch-based methods fundamentally cannot achieve this ideal scenario.
@@
(2) Efficiency issues arising from the fixed and redundant number of tokens limit the application of transformer-based models to remote sensing images. For example, when the patch size is 16 $\times$ 16, a 224 $\times$ 224 remote sensing image corresponds to 196 tokens, and the number of tokens grows quadratically with increasing image size. Furthermore, in transformer models, the computational complexity of the attention mechanism is $O(N^{2})$, which indicates that the computational cost of the attention mechanism also scales quadratically with the number of tokens. This feature makes it challenging for transformer-based remote sensing models to handle high-resolution remote sensing images.
@@
Therefore, meaningless rectangular patches are not suitable as the basic elements of images. Analogous to meaningful words or subwords that serve as the basic elements of natural language, an intuitive idea is to identify certain semantically meaningful regions in the image as the basic elements. Therefore, as illustrated in Figure 1-d, we define semantically independent regions (SIRs) and propose two properties that an ideal visual tokenizer should possess:
(1) Homogeneity: SIRs serve as the basic elements of vision. In this context, an image is a union of all SIRs rather than meaningless rectangular patches.
(2) Adaptivity: The number of tokens should be flexibly adjusted to accommodate images of any size and tasks of varying granularity. Larger images and more fine-grained tasks require more visual tokens, whereas smaller images and simpler tasks need fewer visual tokens.
@@
To develop an ideal visual tokenizer that meets the above properties, we define a confusion matrix for tokens and objects under strict definitions and constraints. We observe that the transformations among ”same object same token”, ”same object multiple tokens”, ”same token multiple objects”, and ”multiple tokens multiple objects” actually involve only two meta-operations: ”split” and ”merge”. The process of obtaining homogeneous visual tokens can be abstracted as a routing selection problem, where two distinct general routes are ”splitting and merging” and ”merging and splitting”.
@@
On the basis of the ”splitting and merging” route, we design a simple HOmogeneous visual tOKenizer: HOOK. HOOK consists of two modules: an object perception module (OPM) and an object vectorization module (OVM).
@@
To achieve homogeneity, HOOK initially splits the image into the finest granularity possible and then gradually merges the split components to form SIRs. In the first step, the OPM uses convolutional blocks to generate 4 $\times$ 4-pixel seeds. The self-attention layers subsequently associate these seeds if they belong to the same SIR. Finally, the OVM employs cross-attention to merge these associated seeds into homogeneous visual tokens.
@@
To achieve adaptivity, the OVM defines $N$ learnable vectors as queries for the cross-attention module, with the seeds output by the OPM serving as the keys and values. Here, $N$ is treated as a variable hyperparameter to allow for the arbitrary adjustment of the token quantity.
@@
We tested the performance of HOOK on a sparse task via the NWPU-RESISC45[18], WHU-RS19[19, 20], and NaSC-TG2[21] classification datasets and on a dense task via the GID5[22] and DGLCC[23] semantic segmentation datasets. The experimental results show that the visual tokens obtained by HOOK correspond to individual objects, thereby verifying their homogeneity. Compared with randomly initialized or pretrained Patch Embed, HOOK achieved performance improvements ranging from 2% to 10%. Additionally, HOOK outperformed the baseline methods that we used for comparison, reaching state-of-the-art performance. In terms of efficiency, HOOK required only 6 tokens for the sparse task and 8 tokens for the dense task, whereas Patch Embed required hundreds of tokens for a single image. HOOK realized an overall efficiency improvement of 1.5 to 2.8 times.
@@
(1) Our analyses and experiments show that the importance of visual tokenizers is far underestimated. Starting from the essence of tokenizers, we propose two fundamental properties that an ideal visual tokenizer should possess: homogeneity and adaptivity.
(2) We define a confusion matrix for tokens and objects to systematically analyse their intricate relationships. This analysis reveals that the process of constructing homogeneous tokens can be abstractly viewed as a routing selection problem involving two principal strategies: ”splitting and merging” and ”merging and splitting”. This discovery provides new perspectives and ideas for research on visual tokenizers.
(3) Based on the ”splitting and merging” strategy, we design a simple homogeneous visual tokenizer, HOOK, which offers an object-oriented, plug-and-play implementation for visual tokenizers. HOOK can obtain visual tokens from semantically independent regions in an image and is adapted to complex objects of any shape rather than just regular rectangular patches.
(4) The experimental results show that HOOK achieves both homogeneity and adaptivity. HOOK outperforms existing baselines in terms of performance and efficiency and has the potential to supplant patch-based methods as a new standard visual tokenizer.
The remainder of this paper is organized as follows: In Section 2, we review the advancements in visual tokenizers, which can be classified into two types, namely, patch-based and object-oriented, based on the different basic elements of vision. In Section 3, we conduct a theoretical analysis of visual tokens. Section 4 and Section 5 introduce HOOK, a homogeneous visual tokenizer designed on the basis of the ”splitting and merging” strategy, and discuss its performance via experiments. Section 6 primarily delves into the significance of visual tokenizers, rethinks our HOOK, and outlines future research directions.
@@ 2 Related Work
2.1 Patch-Based Visual Tokenizers
Patch-based methods use rectangular patches as the basic elements of images. Patch Embed, which is widely applied in patch-based methods, splits an image into nonoverlapping patches of a fixed size, typically 16$\times$16 pixels. The advantage of Patch Embed lies in its simplicity and efficiency. However, the use of redundant tokens and a fixed patch size have been identified in various works[24, 25, 26, 27] as factors resulting in the instability of training in ViT models.
As shown in Figure 1-b, we categorize improved patch-based methods into three types on the basis of hierarchy: patch-level, subpatch-level, and superpatch-level methods.
@@ A. Patch-level Methods
Patch-level methods are based on the argument that the visual tokens obtained through Patch Embed are highly redundant, and in practice, a model needs to select only a small portion of these tokens to recognize an image.
H. Yin et al. proposed A-ViT[28], where the first dimension of each token is defined as the probability of discarding the token, which enables the model to autonomously select tokens that aid in image recognition. Y. Tang et al. calculated the impact of each token at each layer on the final model output, identifying and removing redundant tokens[29]. Y. Rao et al. introduced DynamicViT[30], and B. Pan et al. proposed IA-RED2[31], both of which insert trainable prediction modules in the middle layers of the model for dynamic token selection. D. Marin et al. utilized pooling strategies to discard redundant tokens[32]. Y. Liu et al. introduced PatchDropout[33], a simple method that randomly drops a certain percentage of patches during model training, leading to efficiency improvements without compromising test accuracy. Y. Liang et al. proposed a token reorganization module for identifying and merging tokens with minimal relevance to the CLS token, thereby reducing the number of tokens[34]. Similarly, T. Wang et al. proposed PnP-DETR[35], which initially extracts features from the original image via ResNet as tokens and then identifies and merges background tokens, effectively enhancing model efficiency and object-level perception capabilities.